OpenStudy (anonymous):

Please help. No clue where to start. A person invested $15,000 in stocks and bonds. Her investment in bonds is $3000 more than one third her investment in stocks. How much did she invest in stocks? How much did she invest in bonds?

OpenStudy (igreen):

a person invested $15,000 in stocks and bonds. Her investment in bonds is $3000 more than one third her investment in stocks. How much did she invest in stocks? let her investment in stocks is $x then her investment in bonds=x/3+3000 now x/3+3000+x=15000 or (x+3x)/3=15000-3000=12000 or 4x=12000*3=36000 or x=36000/4=9000 so her investment in stocks is $9000 & in bonds is $6000 http://www.algebra.com/algebra/homework/word/finance/Money_Word_Problems.faq.question.701382.html

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @mjidkc

OpenStudy (anonymous):

Thank you! sorry just saw this!